OpenStudy (anonymous):

Solve the system. x – 2y = 10 3x + 2y = 14 (–8, 6) (–2, 6) (6, –8) (6, –2)

OpenStudy (anonymous):

First, add the two fractions together, since the y values are opposite, they will cancel. You can therefore solve for x first. ( X-y=10) +(3x+2y=14) = (4x = 24) = x = 6. since x = 6, 6 -2y=10 -2y = 4 Y = -2 (6,-2)

OpenStudy (anonymous):

can you help me with this too? @MRamosville Select the pair of equations whose graphs are perpendicular. 5x – 8y = 9 12x – 5y = 7 x + 6y = 8 y = 2x – 8 y = 2x – 7 x + 2y = 3 2y = –3x + 5 2x + 3y = 4